### v2

# Explicit Mutual Information Maximization for Self-Supervised Learning

Abstract
Recently, self-supervised learning (SSL) has been extensively studied. Theoretically, mutual information maximization (MIM) is an optimal criterion for SSL, with a strong theoretical foundation in information theory. However, it is difficult to directly apply MIM in SSL since the data distribution is not analytically available in applications. In practice, many existing methods can be viewed as approximate implementations of the MIM criterion. This work shows that, based on the invariance property of MI, explicit MI maximization can be applied to SSL under a generic distribution assumption, i.e., a relaxed condition of the data distribution. We further illustrate this by analyzing the generalized Gaussian distribution. Based on this result, we derive a loss function based on the MIM criterion using only second-order statistics. We implement the new loss for SSL and demonstrate its effectiveness via extensive experiments.
Index Terms:
self-supervised learning, computer vision, mutual information
I Introduction
Self-supervised learning (SSL) is a powerful technique aims to learn task-agnostic representations without relying on annotated training data. SSL methods commonly utilize pretext tasks to learn representation-extracting models, which can be generally categorized into generative-based and contrastive-based ones [1, 2]. In generative-based methods, the pretext tasks are designed to predict or reconstruct the input data [3, 4, 5]. In comparison, contrastive-based methods construct pretext tasks in the embedding space, and learn a representation-extracting model by discriminating between positive and negative pairs in the embedding space.
Recently, Siamese networks [6] based SSL methods have shown great promise, which utilize weight-sharing networks to process distorted versions of the samples [7, 8, 9, 10, 11]. An issue of Siamese networks is that the two branches may collapse to constant representation. To prevent this problem, contrastive based methods construct positive and negative pairs and contrast between them using contrastive losses, such as InfoNCE loss [12], NT-Xent loss [7], Circle loss [13], and Triplet loss [14]. Besides, online clustering, asymmetric structure and momentum encoder have been shown to be effective for preventing collapsing [10, 15, 16, 11]. Furthermore, redundancy reduction of the embeddings and variance regularization can naturally avoid collapsing [8, 9]. Generally, the primary objective of these methods is to maximize the similarity between the two-branch embeddings of the Siamese networks while avoiding the collapsing problem.
Mutual information (MI) is an information-theoretic measure that can capture non-linear statistical dependencies between random variables and thereby serve as a measure of true dependence [17]. Theoretically, under the framework of Siamese networks based SSL, MI is an optimal metric to measure the dependence between the embeddings of the Siamese networks. However, directly constructing an objective based on MI is challenging as it relies on the analytic expression of the data distribution. Consequently, many existing methods can be seen as approximate implementations of the MI maximization criterion. For example, the InfoNCE loss based methods [7] employ a noise contrastive estimation (NCE) of MI [18], which is shown to be a lower bound of MI [19]. Besides, it has been shown in [20] that the VICReg method [9] implements an approximation of the MI maximization criterion.
In this work, based on the invariance property of MI, we show that explicit MI maximization can be applied to SSL under a generic condition of the data distribution and derive a new objective for SSL. This new objective uses only second-order statistics but involves intractable determinant calculation of high-dimensional matrices. We propose reformulation and optimization strategies to make it adequate for stable and efficient end-to-end training. The main contributions are summarized as follows:
  • We provide a new perspective for SSL objective design from the invariance property of MI established in information theory, which shows that explicit MI optimization can be applied to SSL under a generic condition of the data distribution. Based on this result, an MI objective using only second-order statistics is derived.
  • We implement the second-order statistics based MI objective for SSL, for which optimization strategies have been proposed to make it adequate for end-to-end self-supervised learning on practical tasks.
  • We demonstrate the effectiveness of the proposed method by evaluation on CIFAR-10/100 and ImageNet-100/1K in comparison with state-of-the-art methods.
II Method
We consider a Siamese network architecture that processes two augmented versions of the same input through parallel identical networks [21], as illustrated in Fig. 1.
Figure 1: The proposed objective explicitly measures the MI $I(Z;Z^{\prime})$ between the embeddings $Z$ and $Z^{\prime}$ generated by two identical networks $f_{\omega}(\cdot)$ that are fed transformed versions of the same sample $S$. Maximizing the MI $I(Z;Z^{\prime})$ not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy $H(Z,Z^{\prime})$, but also maximizes their marginal entropy $H(Z)$ and $H(Z^{\prime})$, respectively, which naturally avoids trivial constant solutions.
Denote the two embeddings by $Z=\left[Z_{1},Z_{2},\cdots,Z_{d}\right]^{T}\in\mathbb{R}^{d}$ and $Z^{\prime}=\left[Z_{1}^{\prime},Z_{2}^{\prime},\cdots,Z_{d}^{\prime}\right]^{T}\in\mathbb{R}^{d}$, respectively, which are generated from two distorted views $X$ and $X^{\prime}$ of the same image $S$ by the network $f_{\omega}(\cdot)$ with trainable parameters $\omega$.
II-A The Maximum Mutual Information Criterion
As illustrated in Figure 1, we consider directly maximizing the MI between $Z$ and $Z^{\prime}$, i.e. a maximum mutual information (MMI) criterion as
[…]
The MI can be expressed as $I\left(Z;Z^{\prime}\right)=H(Z)+H(Z^{\prime})-H(Z,Z^{\prime})$, where $H(Z,Z^{\prime})$ is the joint entropy of $(Z,Z^{\prime})$, $H(Z)$ and $H(Z^{\prime})$ are the marginal entropy of $Z$ and $Z^{\prime}$, respectively. It is easy to see that the MMI criterion not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy, but also maximizes the marginal entropy of $Z$ and $Z^{\prime}$, respectively, which promotes decorrelating the features in the embeddings $Z$ and $Z^{\prime}$ and naturally avoids the collapse problem of converging to trivial constant solutions.
However, MI is difficult to explicitly compute except for certain simple specific distributions. The next result shows that $I\left(Z;Z^{\prime}\right)$ can be explicitly computed based only on second-order statistics under a more generic distribution condition. In the following, we denote the covariance matrices of $Z,Z^{\prime}$ and $\tilde{Z}:=\left[Z^{T},Z^{\prime T}\right]^{T}$ by $C_{ZZ}$, $C_{Z^{\prime}Z^{\prime}}$ and $C_{\tilde{Z}\tilde{Z}}$, respectively.
Theorem 1: If there exist two homeomorphisms maps $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ (i.e. $F$ and $G$ are smooth and uniquely invertible maps) such that $Y=F(Z)$ and $Y^{\prime}=G\left(Z^{\prime}\right)$ are Gaussian distributed, let $\tilde{Y}={[{Y^{T}},{Y^{\prime T}}]^{T}}$ and denote the covariance matrices of $Y$, $Y^{\prime}$ and $\tilde{Y}$ by $C_{YY}$, $C_{Y^{\prime}Y^{\prime}}$ and $C_{\tilde{Y}\tilde{Y}}$, respectively, then
$I\left(Z;Z^{\prime}\right)=I\left(Y;Y^{\prime}\right)=\frac{1}{2}\log\frac{\operatorname{det}\left(C_{YY}\right)\operatorname{det}\left(C_{Y^{\prime}Y^{\prime}}\right)}{\operatorname{det}\left(C_{\tilde{Y}\tilde{Y}}\right)}.$ (2)
The proof is given in the Supplementary Material (SM) -B. This result is derived from the invariance property of MI [22] and the MI of multivariate Gaussian distribution. It relaxes the condition of data distribution required for explicit MI expression. It implies that we can compute the MI only based on second-order statistics even if the distributions of $Z$ and $Z^{\prime}$ are not Gaussian. To verify this result, we investigate the MI under a class of distributions, namely generalized Gaussian distribution (GGD). Generally, there does not appear to exist an agree on multivariate GGD, we use a definition as a particular case of the Kotz-type distribution [23, 24]
$\mathcal{GN}\left(X;\mu_{X},\Sigma_{XX},\beta\right):=$ (3)
$\frac{\Phi(\beta,n)}{\left[\operatorname{det}\left(\Sigma_{XX}\right)\right]^{1/2}}\exp\left(-\frac{1}{2}\left[\left(X-\mu_{X}\right)^{T}\Sigma_{XX}^{-1}\left(X-\mu_{X}\right)\right]^{\beta}\right),$
where $\Gamma$ is the gamma function, $\Phi(\beta,n)=\beta\Gamma(n/2)/\left[2^{n/(2\beta)}\pi^{n/2}\Gamma(n/(2\beta)]\right.$, $\mu_{X}$ and $\Sigma_{XX}$ are the mean and dispersion matrix of $X$, respectively, $\beta>0$ is a shape parameter. GGD has a flexible parametric form, which can adapt to a large family of distributions by choosing the shape parameter $\beta$, from super-Gaussian when $\beta<1$ to sub-Gaussian when $\beta>1$, including the Gamma, Laplacian and Gaussian distributions as special cases.
Theorem 2: Suppose that $Z$ and $Z^{\prime}$ follow multivariate GGD, $Z\sim\mathcal{GN}\left(\mu_{Z},\Sigma_{ZZ},\beta\right),Z^{\prime}\sim\mathcal{GN}\left(\mu_{Z^{\prime}},\Sigma_{Z^{\prime}Z^{\prime}},\beta\right)$, where $\Sigma_{ZZ}$ and $\Sigma_{Z^{\prime}Z^{\prime}}$ are dispersion matrices. Let $\tilde{Z}=\left[Z^{T},Z^{\prime T}\right]^{T}\in\mathbb{R}^{2d}$ and denote the dispersion matrix of $\tilde{Z}$ by $\Sigma_{\tilde{Z}\tilde{Z}}$. Then, for any $\beta>0$, the MI between $Z$ and $Z^{\prime}$ is given by
$I\left(Z;Z^{\prime}\right)$ $=\frac{1}{2}\log\frac{\operatorname{det}\left(\Sigma_{ZZ}\right)\operatorname{det}\left(\Sigma_{Z^{\prime}Z^{\prime}}\right)}{\operatorname{det}\left(\Sigma_{\tilde{Z}\tilde{Z}}\right)}$ (4)
$=\frac{1}{2}\log\frac{\operatorname{det}\left(C_{ZZ}\right)\operatorname{det}\left(C_{Z^{\prime}{Z}^{\prime}}\right)}{\operatorname{det}\left(C_{\tilde{Z}\tilde{Z}}\right)}.$
The proof is given in SM -C. The above result implies that MI can be explicitly computed based only on second-order statistics for any distribution that can be marginally reparameterized as multivariate norm distribution under a homeomorphism condition.
Based on the above results, we propose to use the second-order statistics based explicit MI expression for SSL, with which the MMI criterion (1) can be formulated as
[…]
Minimizing formulation (5) is equivalent to maximizing $\operatorname{det}\left(C_{ZZ}\right)$ and $\operatorname{det}\left(C_{Z^{\prime}Z^{\prime}}\right)$ while minimizing $\operatorname{det}\left(C_{\tilde{Z}\tilde{Z}}\right)$. Maximizing $\operatorname{det}\left(C_{ZZ}\right)$ (resp. $\operatorname{det}\left(C_{Z^{\prime}Z^{\prime}}\right)$ ) would promote the features in $Z$ (resp. $Z^{\prime}$ ) to be less correlated, therefore decorrelates the features in the output embedding. This decorrelation reduces the redundant information of the output embedding and promotes extracting information from the sample as much as possible, which also helps avoid the collapse problem of converging to a trivial constant solution.
II-B Implementation of the MMI Criterion for SSL
It is difficult to directly optimize the MMI formulation (5) as it involves computing the determinant of high-dimensional covariance matrices, e.g., $C_{ZZ}\in\mathbb{R}^{d\times d},C_{Z^{\prime}Z^{\prime}}\in\mathbb{R}^{d\times d}$, and $C_{\tilde{Z}\tilde{Z}}\in\mathbb{R}^{2d\times 2d}$. Typically, $m<d$ in the self-supervised learning setting, e.g., $d=4096,m\in\{512,1024,2048\}$. The MMI formulation can be reformulated to reduce the computational complexity. Specifically, since the two transformation operators $X=t\left(S\right)$ and $X^{\prime}=t^{\prime}\left(S\right)$ are sampled from the same distribution of augmentations, i.e. $t\sim T$ and $t^{\prime}\sim T,Z=f_{\omega}(X)$ and $Z^{\prime}=f_{\omega}\left(X^{\prime}\right)$ are the representations extracted by the same network $f_{\omega}$, and without loss of generality assuming that $Z$ and $Z^{\prime}$ having zero-mean, it follows that
$C_{ZZ}=E\left\{ZZ^{T}\right\}=C_{Z^{\prime}Z^{\prime}}=E\left\{Z^{\prime}Z^{\prime T}\right\},$
$C_{ZZ^{\prime}}=E\left\{ZZ^{\prime T}\right\}=C_{Z^{\prime}Z}=E\left\{Z^{\prime}Z^{T}\right\}.$
Then, using $\operatorname{det}\left(\left[\begin{array}[]{ll}\!A&B\!\\ \!B&A\!\end{array}\right]\right)=\operatorname{det}(A+B)\operatorname{det}(A-B)$, we have
$\operatorname{det}\left(C_{\tilde{Z}\tilde{Z}}\right)$ $=\operatorname{det}\left(\left[\begin{array}[]{cc}\!C_{ZZ}&C_{ZZ^{\prime}}\!\\ \!C_{Z^{\prime}Z}&C_{Z^{\prime}Z^{\prime}}\!\end{array}\right]\right)$
$=\operatorname{det}\left(C_{ZZ}+C_{ZZ^{\prime}}\right)\operatorname{det}\left(C_{ZZ}-C_{ZZ^{\prime}}\right).$
Therefore, the objective of the MMI formulation (5) can be reformulated as
$\mathcal{L}$ $=\log\operatorname{det}\left(C_{ZZ}+C_{ZZ^{\prime}}\right)+\log\operatorname{det}\left(C_{ZZ}-C_{ZZ^{\prime}}\right)$ (6)
$-\log\operatorname{det}\left(C_{ZZ}\right)-\log\operatorname{det}\left(C_{Z^{\prime}Z^{\prime}}\right).$
In the following without loss of generality we assume that $Z$ and $Z^{\prime}$ are normalized such that $\operatorname{diag}\left\{C_{ZZ}\right\}=\operatorname{diag}\left\{C_{Z^{\prime}Z^{\prime}}\right\}=I$. Then, it is easy to see that, the desired optimal solution of minimizing (6) is $C_{ZZ}=\pm C_{ZZ^{\prime}}$ and $C_{ZZ}=C_{Z^{\prime}Z^{\prime}}=I$. That is $Z=\pm Z^{\prime}$, and, at meantime, $Z$ and $Z^{\prime}$ have the maximum entropy given by $C_{ZZ}=C_{Z^{\prime}Z^{\prime}}=I$. However, from the above analysis, since $X=t\left(S\right)$ and $X^{\prime}=t^{\prime}\left(S\right)$ with $t(\cdot)$ and $t^{\prime}(\cdot)$ following the same distribution, $Z=f_{\omega}(X)$ and $Z^{\prime}=f_{\omega}\left(X^{\prime}\right)$ are the outputs of the same network $f_{\omega}$, the solution $C_{ZZ}=-C_{ZZ^{\prime}}$ given by $Z=-Z^{\prime}$ is invalid. Hence, we can drop the first term in (6) to obtain
$\mathcal{L}=\log\operatorname{det}\left(C_{ZZ}-C_{ZZ^{\prime}}\right)-\log\operatorname{det}\left(C_{ZZ}\right)-\log\operatorname{det}\left(C_{Z^{\prime}Z^{\prime}}\right).$ (7)
It is problematic to directly optimize the objective (7). First, the MI is formulated for multivariate continuous variables, which would be infinite when any two components are linearly correlated, e.g., $I\left(Z;Z^{\prime}\right)=\infty$ if $Z_{i}=\phi Z_{i}^{\prime}$ for any $i\in\{1,\cdots,d\}$ and any $\phi\neq 0$. This means $\mathcal{L}=-\infty$ when any pair $\left(Z_{i},Z_{i}^{\prime}\right)$ is linearly correlated, regardless the correlation between other feature pairs $\left\{\left(Z_{j},Z_{j}^{\prime}\right)\right\}_{j\neq i}$. Second, the high-dimensional determinant is intractable to compute, and direct minimization of (7) has a degeneration problem when $m<d$ since the covariance matrices are rank-deficient in this case. To achieve stable optimization for end-to-end training, we consider further reformulation and approximation of (7) as follows. Specifically, in implementation the covariance matrices are estimated from $m$ samples, and an empirical loss of (7) is used as
$\mathcal{L}=\log\operatorname{det}(\hat{C}_{ZZ}-\hat{C}_{ZZ^{\prime}})-\log\operatorname{det}(\hat{C}_{ZZ})-\log\operatorname{det}(\hat{C}_{Z^{\prime}Z^{\prime}}).$ (8)
With $m<d$, the matrices $\hat{C}_{ZZ},\hat{C}_{ZZ^{\prime}},\hat{C}_{Z^{\prime}Z^{\prime}}$ are rank-deficient and their determinants are constantly zero, which makes (8) an inadequate objective for optimizing the network. In this case, recalling that the determinant of a matrix is equivalent to the product of its eigenvalues, we turn to optimize the product of its nonzero eigenvalues. Under the assumption that $Z$ and $Z^{\prime}$ have zero-mean, and let $\bar{Z}\in\mathbb{R}^{d\times m}$ and $\bar{Z}^{\prime}\in\mathbb{R}^{d\times m}$ denote the matrices containing the embeddings of $m$ samples, this can be implemented by replacing $\hat{C}_{ZZ}=\frac{1}{d}\bar{Z}\bar{Z}^{T},\hat{C}_{ZZ^{\prime}}=\frac{1}{d}\bar{Z}\bar{Z}^{\prime T},\hat{C}_{Z^{\prime}Z^{\prime}}=\frac{1}{d}\bar{Z}^{\prime}\bar{Z}^{\prime T}$ by $\breve{C}_{ZZ}=\frac{1}{m}\bar{Z}^{T}\bar{Z},\breve{C}_{ZZ^{\prime}}=\frac{1}{m}\bar{Z}^{T}\bar{Z}^{\prime}$, and $\breve{C}_{Z^{\prime}Z^{\prime}}=\frac{1}{m}\bar{Z}^{\prime T}\bar{Z}^{\prime}$, respectively, to get
$\mathcal{L}\!=\!\log\operatorname{det}\left(\bar{Z}^{T}\bar{Z}-\bar{Z}^{T}\bar{Z}^{\prime}\right)-\log\operatorname{det}\left(\bar{Z}^{T}\bar{Z}\right)-\log\operatorname{det}\left(\bar{Z}^{\prime T}\bar{Z}^{\prime}\right).$ (9)
Besides, since the determinant of a high dimensional matrix is expensive to compute and unstable, we consider a Tylor expansion of the log-determinant $\log\operatorname{det}(M)$ for a matrix $M\in\mathbb{R}^{n\times n}$ as
$\log\operatorname{det}(M)=\sum_{i=1}^{n}\log\lambda_{i}(M)$
$=\sum_{i=1}^{n}\sum_{k=1}^{\infty}(-1)^{k+1}\frac{\left(\lambda_{i}(M)-1\right)^{k}}{k}$
$=\sum_{k=1}^{\infty}(-1)^{k+1}\frac{\operatorname{tr}\left((M-I)^{k}\right)}{k}\operatorname{tr}\left(\sum_{k=1}^{\infty}(-1)^{k+1}\frac{(M-I)^{k}}{k}\right),$ (10)
where we used $\sum_{i=1}^{n}\left(\lambda_{i}(M)-1\right)^{k}=\operatorname{tr}\left((M-I)^{k}\right)$, and $\lambda_{i}(M)$ denotes the $i$-th eigenvalue of $M$. The series (-E) only converges under the condition of $\|M-I\|<1$, or equivalently $20<\lambda_{i}(M)<$ 2 for any $i$. We seek to use a low-order approximation of (-E) to facilitate the optimization. From (-E), since a lower-order approximation is more accurate when the eigenvalues $\left\{\lambda_{i}(M)\right\}_{i=1,\ldots,n}$ are closer to 1, we rescale the matrix as
$\tilde{M}=\frac{M-\mu_{\lambda}I}{\alpha}+I,$ (11)
where $\mu_{\lambda}=\frac{\lambda_{\max}(M)+\lambda_{\min}(M)}{2}$ and $\alpha=\beta\left(\mu_{\lambda}-\lambda_{\text{min }}(M)\right)$, such that $\|\tilde{M}-I\|\leq\frac{1}{\beta}<1$ for some $\beta>1.~{}\lambda_{\text{min }}(\cdot)$ and $\lambda_{\text{max }}(\cdot)$ denote the minimum and maximum eigenvalues of a matrix, respectively. In implementing the loss (9), the three log-determinant terms are expanded as (-E), and only a $p$-th order approximation is kept, e.g., $p=4$ in the experiments of this work. A fourth-order approximation of the log function in (-E) is sufficiently accurate around the value of 1.
Notice that using the diagonal-loading like rescaling in (11) is equivalent to using a noise injection regularization of the MI calculation. Specifically, it is equivalent to replacing $I\left(Z;Z^{\prime}\right)$ by $I\left(Z+\varepsilon;Z^{\prime}+\varepsilon^{\prime}\right)$, where $\varepsilon$ and $\varepsilon^{\prime}$ are additive Gaussian noise. With this regularization the MI becomes finite. Otherwise, the MI would be infinite when any two components are linearly correlated, e.g., $I\left(Z;Z^{\prime}\right)=\infty$ if $Z_{i}=\phi Z_{i}^{\prime}$ for any $i\in\{1,\cdots,d\}$ and any $\phi\neq 0$, and correspondingly with the first term of the objective (5) becoming minus infinite, which makes the objective in (5) difficult to optimize directly.
III Experiments
We follow the standard training protocol in the Solo-learn benchmark [25]. ResNet18 is used as the backbone for CIFAR-10/100 and ImageNet-100, whilst ResNet50 is used for ImageNet-1K. The results on CIFAR-10/100 and ImageNet-100 in Table I are compared with a batch size of 256. The experiments are conducted on a computing platform with three A100 GPUs, each with 40 GB of memory. On ImageNet-1K, it supports a maximum batch size of 1020 when using 16-bit precision (FP16). Therefore, on ImageNet-1K, our method is run with a precision of FP16 and a batch size of 1020. Meanwhile, to enhance stability and expedite the training process, we use gradient accumulation across four batches ($1020\times 4$) on ImageNet-1K. Detailed experiment settings are provided in the SM -E.
TABLE I: Linear probing results on CIFAR-10/100 and ImageNet-100. Top-1 accuracy is reported with ResNet18.
Method CIFAR-10 CIFAR-100 ImageNet-100
Barlow Twins [8] 92.1 70.9 80.4
BYOL[15] 92.6 70.5 80.2
DeepCluster V2 [10] 88.9 63.6 75.4
DINO [26] 89.5 66.8 74.8
MoCo V2+ [27] 92.9 69.9 78.2
MoCo V3 [28] 93.1 68.8 80.4
NNCLR [29] 91.9 69.6 79.8
ReSSL [30] 90.6 65.9 76.9
SimCLR [7] 90.7 65.8 77.6
Simsiam [16] 90.5 66.0 74.5
SwAV [31] 89.2 64.9 74.0
VICReg [9] 92.1 68.5 79.2
W-MSE [32] 88.7 61.3 67.6
Ours 93.1 70.5 81.1
TABLE II: Linear probing results on ImageNet-1K. Top-1 accuracy is reported at 100, 400, and 800 epochs with ResNet50.
Method 100 eps 400 eps 800 eps
SimCLR [7] 66.5 69.8 70.4
MoCo v2 [27] 67.4 71.0 72.2
BYOL [15] 66.5 73.2 74.3
SwAV [31] 66.5 70.7 71.8
SimSiam [16] 68.1 70.8 71.3
Barlow Twins [8] 67.3 71.8 73.0
Ours 70.6 72.6 73.1
LINEAR PROBING
Linear probing is a standard evaluation protocol in SSL, in which a linear classifier is trained on top of frozen representations of a pretrained model. Table I shows the linear probing results on CIFAR-10/100 and ImageNet-100. Following strictly the protocol of the Solo-learn benchmark [25], ResNet18 is used with the last fully connected layer replaced with a three-layer MLP interleaved with Batch Normalization (BN) and ReLU as projector. We keep the projector’s hidden dimension and output dimension as 2048. The modified backbone and projector act as encoder. For CIFAR, we adapt the architecture by removing the first max-pooling layer and modifying the first convolutional layer to fit the $32\times 32$ input size, in line with [25, 33, 7]. The results for other methods in Table I are sourced from the Solo-learn benchmark [25]. It can be seen that our method achieves competitive performance on all three datasets. Notably, on ImageNet-100, our method achieves the highest accuracy of 81.1%. Table II compares the top-1 accuracy of the methods on ImageNet-1K with different pretraining epochs. On ImageNet-1K, the results of our method are reported with a momentum encoder [15, 16]. As will be shown in Table III, our method can achieve better performance with a momentum encoder. It can be seen from Table II that our method can achieve competitive performance on ImageNet-1K, which is especially noteworthy under small training epochs.
ABLATION ON BATCH SIZE
We evaluate the robustness of our method against batch size, following the same setting as Table I. We consider two variants of our method, one is standard and the other uses a momentum encoder [15, 16]. As shown in Table III, both our method and Barlow Twins reach a saturation point as batch size increases, which is consistent with the results in [8, 33]. Our method has better robustness across varying batch sizes, with a maximum accuracy variation of 1.6% on CIFAR-100 and 2.0% on ImageNet-100, respectively, while Barlow Twins shows 5.6% and 2.9% accuracy variation on the two datasets. For our method, using a momentum encoder helps to enhance performance across different batch sizes. More ablation studies can refer to the SM -D.
TABLE III: Robustness to batch size. We use the same setup as in Table I but vary the batch size. “Ours-M” denotes our method with a momentum encoder.
Batch size CIFAR-100 ImageNet-100
Ours Ours-M Barlow Twins Ours Ours-M Barlow Twins
1024 68.9 70.4 65.9 79.8 81.4 77.5
512 70.0 70.5 68.7 79.5 81.1 79.6
256 70.5 70.4 70.9 81.1 81.7 80.4
128 70.2 70.2 71.5 80.7 81.8 80.2
64 70.1 71.1 70.2 79.1 81.2 78.4
IV Conclusions
This work seeks to construct an objective for SSL from the information theoretic view. Based on the invariance property of mutual information, we showed that the maximum MI criterion can be applied to SSL for a relaxed condition on the data distribution. We illustrated this through analyzing the explicit MI of generalized Gaussian distribution. Then, the derived second-order statistics based formulation of MI is employed for SSL, and optimization strategies have been proposed to efficiently implement it for end-to-end training. Experiments on the CIFAR and ImageNet benchmarks demonstrated the effectiveness of the derived MI based loss.

### v3

@@
 SSL methods commonly utilize pretext tasks to learn representation-extracting models, which can be generally categorized into generative-based and contrastive-based ones [1, 2].
-In generative-based methods, the pretext tasks are designed to predict or reconstruct the input data [3, 4, 5].
+In generative-based methods, the pretext tasks are designed to predict or reconstruct the input data [3, 4].
 In comparison, contrastive-based methods construct pretext tasks in the embedding space, and learn a representation-extracting model by discriminating between positive and negative pairs in the embedding space.
@@
 Besides, it has been shown in [19] that the VICReg method [8] implements an approximation of the MI maximization criterion.
-In this work, based on the invariance property of MI, we show that explicit MI maximization can be applied to SSL under a generic condition of the data distribution and derive a new objective for SSL.
-This new objective uses only second-order statistics but involves intractable determinant calculation of high-dimensional matrices.
-We propose reformulation and optimization strategies to make it adequate for stable and efficient end-to-end training.
-The main contributions are summarized as follows:
-  • We provide a new perspective for SSL objective design from the invariance property of MI established in information theory, which shows that explicit MI optimization can be applied to SSL under a generic condition of the data distribution.
-Based on this result, an MI objective using only second-order statistics is derived.
+In [20], the minimal coding length in lossy coding is used to construct a maximum entropy coding objective for SSL.
+Meanwhile, MI maximization is considered in [21] for SSL using a log-determinant approximation of the MI, which is simplified to derive a Euclidean distance-based objective.
+In this work, based on the invariance property of MI, we show that the explicit MI maximization objective based on second-order statistics can be applied to SSL under a generic condition of the data distribution.
+This objective involves intractable determinant calculation of high-dimensional matrices, we propose reformulation and optimization strategies to make it adequate for stable and efficient end-to-end training.
+The main contributions are as follows:
+  • We provide a new perspective for SSL objective design from the invariance property of MI established in information theory, which shows that explicit MI optimization based on second-order statistics can be applied to SSL under a generic condition of the data distribution.
   • We implement the second-order statistics based MI objective for SSL, for which optimization strategies have been proposed to make it adequate for end-to-end self-supervised learning on practical tasks.
@@
 We consider a Siamese network architecture that processes two augmented versions of the same input through parallel identical networks [22], as illustrated in Fig. 1.
-Figure 1: The proposed objective explicitly measures the MI $I(Z;Z^{\prime})$ between the embeddings $Z$ and $Z^{\prime}$ generated by two identical networks $f_{\omega}(\cdot)$ that are fed transformed versions of the same sample $S$.
-Maximizing the MI $I(Z;Z^{\prime})$ not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy $H(Z,Z^{\prime})$, but also maximizes their marginal entropy $H(Z)$ and $H(Z^{\prime})$, respectively, which naturally avoids trivial constant solutions.
+Figure 1: The MMI objective explicitly measures the MI $I(Z;Z^{\prime})$ between the embeddings $Z$ and $Z^{\prime}$ generated by two identical networks $f_{\omega}(\cdot)$ that are fed transformed versions of sample $S$.
+Maximizing $I(Z;Z^{\prime})$ not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy $H(Z,Z^{\prime})$, but also maximizes their marginal entropy $H(Z)$ and $H(Z^{\prime})$, respectively, which naturally avoids trivial constant solutions.
 Denote the two embeddings by $Z=\left[Z_{1},Z_{2},\cdots,Z_{d}\right]^{T}\in\mathbb{R}^{d}$ and $Z^{\prime}=\left[Z_{1}^{\prime},Z_{2}^{\prime},\cdots,Z_{d}^{\prime}\right]^{T}\in\mathbb{R}^{d}$, respectively, which are generated from two distorted views $X$ and $X^{\prime}$ of the same image $S$ by the network $f_{\omega}(\cdot)$ with trainable parameters $\omega$.
 II-A The Maximum Mutual Information Criterion
-As illustrated in Figure 1, we consider directly maximizing the MI between $Z$ and $Z^{\prime}$, i.e. a maximum mutual information (MMI) criterion as
+As shown in Fig. 1, we consider a maximum mutual information (MMI) criterion as
 […]
 The MI can be expressed as $I\left(Z;Z^{\prime}\right)=H(Z)+H(Z^{\prime})-H(Z,Z^{\prime})$, where $H(Z,Z^{\prime})$ is the joint entropy of $(Z,Z^{\prime})$, $H(Z)$ and $H(Z^{\prime})$ are the marginal entropy of $Z$ and $Z^{\prime}$, respectively.
-It is easy to see that the MMI criterion not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy, but also maximizes the marginal entropy of $Z$ and $Z^{\prime}$, respectively, which promotes decorrelating the features in the embeddings $Z$ and $Z^{\prime}$ and naturally avoids the collapse problem of converging to trivial constant solutions.
+The MMI criterion not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy, but also maximizes the marginal entropy of $Z$ and $Z^{\prime}$, respectively, which promotes decorrelating the features in the embeddings $Z$ and $Z^{\prime}$ and naturally avoids the collapse problem of converging to trivial constant solutions.
 However, MI is difficult to explicitly compute except for certain simple specific distributions.
+Accordingly, Gaussian distribution assumption is used to derive MI based objectives for SSL [21, 19].
 The next result shows that $I\left(Z;Z^{\prime}\right)$ can be explicitly computed based only on second-order statistics under a more generic distribution condition.
@@
 $I\left(Z;Z^{\prime}\right)=I\left(Y;Y^{\prime}\right)=\frac{1}{2}\log\frac{\operatorname{det}\left(C_{YY}\right)\operatorname{det}\left(C_{Y^{\prime}Y^{\prime}}\right)}{\operatorname{det}\left(C_{\tilde{Y}\tilde{Y}}\right)}.$ (2)
-The proof is given in the Supplementary Material (SM) -B.
+The proof is given in the supplemental material (SM) [23].
 This result is derived from the invariance property of MI [24] and the MI of multivariate Gaussian distribution.
@@
 $=\frac{1}{2}\log\frac{\operatorname{det}\left(C_{ZZ}\right)\operatorname{det}\left(C_{Z^{\prime}{Z}^{\prime}}\right)}{\operatorname{det}\left(C_{\tilde{Z}\tilde{Z}}\right)}.$
-The proof is given in SM -C.
+The proof is given in the SM [23].
 The above result implies that MI can be explicitly computed based only on second-order statistics for any distribution that can be marginally reparameterized as multivariate norm distribution under a homeomorphism condition.
@@
 III Experiments
-We follow the standard training protocol in the Solo-learn benchmark [25].
+We follow the standard training protocol in Solo-learn benchmark [27].
 ResNet18 is used as the backbone for CIFAR-10/100 and ImageNet-100, whilst ResNet50 is used for ImageNet-1K.
@@
 Meanwhile, to enhance stability and expedite the training process, we use gradient accumulation across four batches ($1020\times 4$) on ImageNet-1K.
-Detailed experiment settings are provided in the SM -E.
+Detailed experiment settings are provided in the SM [23].
 TABLE I: Linear probing results on CIFAR-10/100 and ImageNet-100. Top-1 accuracy is reported with ResNet18.
@@
 W-MSE [34] 88.7 61.3 67.6
+CorInfoMax [21] 93.2 71.6 80.5
 Ours 93.1 70.5 81.1
@@
 The modified backbone and projector act as encoder.
-For CIFAR, we adapt the architecture by removing the first max-pooling layer and modifying the first convolutional layer to fit the $32\times 32$ input size, in line with [25, 33, 7].
+For CIFAR, we adapt the architecture by removing the first max-pooling layer and modifying the first convolutional layer to fit the $32\times 32$ input size, in line with [27, 6].
 The results for other methods in Table I are sourced from the Solo-learn benchmark [27].
@@
 On ImageNet-1K, the results of our method are reported with a momentum encoder [14, 15].
-As will be shown in Table III, our method can achieve better performance with a momentum encoder.
-It can be seen from Table II that our method can achieve competitive performance on ImageNet-1K, which is especially noteworthy under small training epochs.
+As shown in Table III, our method achieves better performance with a momentum encoder.
+From Table II, our method achieves competitive performance on ImageNet-1K, which is especially noteworthy under small training epochs.
 ABLATION ON BATCH SIZE
@@
 We consider two variants of our method, one is standard and the other uses a momentum encoder [14, 15].
-As shown in Table III, both our method and Barlow Twins reach a saturation point as batch size increases, which is consistent with the results in [8, 33].
+As shown in Table III, both our method and Barlow Twins reach a saturation point as batch size increases, which is consistent with the results in [7].
 Our method has better robustness across varying batch sizes, with a maximum accuracy variation of 1.6% on CIFAR-100 and 2.0% on ImageNet-100, respectively, while Barlow Twins shows 5.6% and 2.9% accuracy variation on the two datasets.
 For our method, using a momentum encoder helps to enhance performance across different batch sizes.
-More ablation studies can refer to the SM -D.
+More ablation studies can refer to the SM [23].
 TABLE III: Robustness to batch size. We use the same setup as in Table I but vary the batch size. “Ours-M” denotes our method with a momentum encoder.